OpenStudy (anonymous):

and how to find the limit x-> a (x*ln(A)-a*ln(B)), if limit ln(A) & ln(B) approach to infinity???????

OpenStudy (zarkon):

\[\ln(A)-\ln(B)=\ln\left(\frac{A}{B}\right)\]

OpenStudy (zarkon):

then try taking the limit